Grade 12Mechanics

a projectile is thrown from base of an incline of angle 30. it is thrown at an angle of 60 from the horizontal direction at a speed 10 m/s . the total time of flight is

arav
8 Years agoGrade 12

1 Answer

swarup kumar majee
8 Years ago
the formula of time of flight on a inclined plane is=2usin\theta/gcos\theta
                                                                             =2*10*1/2\div9.8*31/2/2
                                                                             =10/8.5=1.17
therefore the total time of flight is 1.17second